OpenStudy (anonymous):

given f(x)=4x^2-3 and g(x)=-3x^2+4,find (f.g)(x) AWAAAAARDDDD!!!1

OpenStudy (anonymous):

\[f\circ g\] or \[f\times g\]?

OpenStudy (anonymous):

F x g @satellite73

OpenStudy (anonymous):

put a times sign between them use parentheses

OpenStudy (anonymous):

what do you mean like (fxg)?? @satellite73

OpenStudy (anonymous):

can u demonstrate because i have 7 problems like this and i want to get a good grade @satellite73

OpenStudy (solomonzelman):

what is ` (4x²-3) × (-3x²+4) ` ?

OpenStudy (solomonzelman):

the first part, the `4x²-3` is f, and the second part, the `-3x²+4 `, is the g. that is what they mean by f × g

OpenStudy (solomonzelman):

all I am trying to say that in the following case, \(\large\color{black}{ \displaystyle f(x)=4x^2-3 ~~~~{\rm and}~~~~ g(x)=-3x^2+42}\) the (f×g)(x) is going to be \(\large\color{black}{ \displaystyle (f×g)(x)=\left(4x^2-3\right)\times \left(-3x^2+42\right)}\)

OpenStudy (solomonzelman):

so you are to multiply what f(x) is equal to, BY what g(x) is equal to.

OpenStudy (solomonzelman):

is this making some sense?

OpenStudy (anonymous):

yes it is ! so do i basically simplify it? @SolomonZelman

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

when you expand that product, tell me what you get:)

OpenStudy (anonymous):

would it be something like 4x^2+2x-3? pr 6x^2-3 or just 4x^3-2 ?????? @SolomonZelman

OpenStudy (solomonzelman):

you are multiplying....

OpenStudy (solomonzelman):

(4x²−3) × (−3x²+42)

OpenStudy (anonymous):

idk this topic sir , its really confusing @SolomonZelman

OpenStudy (anonymous):

for this question : 

OpenStudy (anonymous):

@SolomonZelman ^^^^

OpenStudy (solomonzelman):

8x³-12x² it would be for the problem you posted just now

OpenStudy (solomonzelman):

i suggest to go back and review multiplication of variables with different coefficients and powers, factoring and all that.